OpenStudy (anonymous):

How do you do this The sum of the digits of a two digit number is one third of the number. the units digit is 5 more than the tens digit. What is the number?

OpenStudy (anonymous):

Let the digits be t (for tens) and o (for ones). t + o = (10t + o)/3 o = t + 5 Substitute and figure out t. Then, figure out o. Then, the number.

OpenStudy (anonymous):

I dont get what to do in the 10t+o/3

OpenStudy (anonymous):

If the tens digit is t and ones digit is o then the number has to be 10*t + o.

OpenStudy (anonymous):

Don't you have to get rid of the 3?

OpenStudy (anonymous):

t + o = (10t+o)/3 => 3t + 3o = 10t + o => 7t -2o = 0 => 7t = 2o But, since o = t+5, you can substitute. => 7t = 2(t+5) => 7t = 2t + 10 => 5t = 10 => t = 2 => o = t +5 => o = 7

OpenStudy (anonymous):

So the answer is 27?

OpenStudy (anonymous):

Yepsie doo!

OpenStudy (anonymous):

THanks!